# An Effective Version of the $p$-Curvature Conjecture for Order One Differential Equations

### v1

@@ 5.5 Complexity Estimate
Let us justify that the $\delta$ from Theorem 1.3 is negligible in our estimates.
@@
Let $\Delta\in\mathbb{N}$ and $\delta=\prod_{p|\Delta}p^{1/(p-1)}$, then $\delta=O(\log(\Delta)^{2})$.
Доказательство.
First we remark that if $\ell\in\mathbb{N}$ is such that $\ell\#\leq\Delta<(\ell+1)\#$, then $\delta(\Delta)\leq\delta(\ell\#)$ where $\ell\#\coloneqq\prod_{p\leq\ell}p$ is the primorial of $\ell$. This follows from the fact that $x^{1/(x-1)}$ is decreasing for $x\geq 2$. We will now estimate $\delta(\ell\#)$. In [RS62, Theorem 10], we find that for $\ell\geq 557$, we have $0.92\cdot\ell\leq\log(\ell\#)$. Checking numerically that for $1\leq\ell\leq 557$ we have $2^{\ell/2}\leq\ell\#$ we conclude that this inequality holds for all $\ell\geq 1$. Moreover we have the estimate $\sum_{p\leq\ell}\frac{\log(p)}{p-1}\leq 2\sum_{p\leq\ell}\frac{\log(p)}{p}\leq 2\log(\ell)$ following [RS62, Eq. (2.5), (2.11)]. Taking now $\ell$ such that $\ell\#\leq\Delta\leq(\ell+1)\#$ we obtain that $\delta(\Delta)\leq\frac{4}{\log(2)^{2}}\log(\Delta)^{2}=O(\log(\Delta)^{2})$. ∎
@@
Given $u(x)\coloneqq\frac{a(x)}{b(x)}\in\mathbb{Q}(x)$ of degree $n$ and multiplicative height $H$, deciding if all solutions of $y^{\prime}(x)=u(x)y(x)$ are algebraic can be done using Algorithm 3 by performing $\tilde{O}(\Delta^{6}B)=\tilde{O}(H^{12n-6}n^{12n}3^{-3n})$ bit operations, where $\tilde{O}$ hides factors polynomial in $n$ and logarithmic in $H$.
Доказательство.
We compute $R(w),\Delta,$ and $\delta$ as described in the algorithm, and the maximal modulus of a root of $R(w)$, up to a constant factor by for instance setting $\tau=\frac{1}{2}$ in Proposition 5.11. This we will use as an upper bound on $B$. With this we obtain an upper bound on $\sigma$. Clearly the complexities of these computations are negligible compared to the computation of sufficiently many $p$-curvatures as described below.
By Theorem 1.3, it suffices to compute the $p$-curvatures up to $\sigma$, defined in its statement. Thus, we want to set $S=\sigma$ in Proposition 5.12. In accordance with Lemma 5.14 we have $\sigma=\tilde{O}(B\Delta^{6})$ where we omit logarithmic factors in $\Delta$. Cauchy’s bound on the roots of $R(w)=r_{n}w^{n}+\dots+r_{0}$ states that we can take $B$ at most $1+\max_{i}\left|\frac{r_{i}}{r_{n}}\right|$, hence $\sigma=\tilde{O}(\Delta^{5}H_{R})$, where $H_{R}$ denotes the multiplicative height of $R(w)$. Using estimates from Proposition 5.3 to estimate this height and $\Delta$ we find both $\Delta$ and $H_{R}$ to be in $\tilde{O}(H^{2n-1}n^{2n-1/2}3^{-n/2})$. Taking the $6$-th power of this quantity and combining with Proposition 5.12 concludes. ∎
@@ 6.1 Algorithmic Improvements
The crucial algorithmic idea of the design of Algorithm 3 is to delay the moment when costly computations are done, that is the computation of the Rothstein-Trager resultant $R(w)$, when the degree of the input increases. By checking the ‘‘first’’ $p$-curvatures, in our implementation for primes $p$ up to $\Delta$, we try to return ‘‘Transcendental’’ before having to compute $R(w)$. In case no $p$-curvature vanishes until this point, the computation of further $p$-curvatures, potentially up to $\sigma$, is inevitable and this has much higher complexity than the computation of $R(w)$.
When computing the bound $B$ on the modulus of the roots of $R(w)$, the precision $\tau$ of the computation influences the complexity, see Proposition 5.11. In our case, computing all $p$-curvatures up to $\sigma$ already exceeds the capabilities of modern computers in most cases. For theoretical estimates, having $B$ multiplied by a constant factor does not change the complexity of the computations. Our implementation relies on native SageMath functions numerically computing zeroes of a polynomial, up to a fixed precision of $2^{-53}$.
@@ 6.2.1 Computation of $p$-curvatures
First of all we would like to ensure and to convince the reader that our implementation of Algorithm 4 of [BS09] to compute $p$-curvatures of first-order differential equations is indeed faster than Pagиs’ implementation of his algorithm [Pag21] applied to order one equations. The following timings in average were noted for both algorithms when asked to compute $p$-curvatures for primes up to $S$ on random polynomial inputs $a(x),b(x)\in\mathbb{Z}[x]$ of given multiplicative height $H$ and degrees n1n-1, respectively $n$.
@@ 33 $2^{80}$ $100$ $25$ ms $140$ ms
Таблица 1: Computation time of two algorithms computing $p$-curvatures.
@@ 6.2.2 Maple’s istranscendental
The maple package gfun [SZ94] has a command called istranscendental based on the algorithms described in [BSS25], aiming to prove transcendence of a given solution of a given differential equation. Among other things, it checks whether the singularities of the minimal differential operator annihilating a given function are all regular, and if its local exponents are rational. This is a necessary criterion for the algebraicity of the solution, so in case one of these conditions is violated, the command outputs true. For order one differential equations however, Proposition 3.3 ensures that the conditions are also sufficient for algebraicity. While istranscendental outputs FAIL in this case, we can actually conclude algebraicity and thus we are (ab)using the command as an implementation of Algorithm 2.
@@ 6.3.1 Small Inputs
$\cfrac{3x-4}{2x^{2}-6x+4}$ $265$ A $120$ ms $45$ ms <1<1 ms $25$ ms
$\cfrac{7x^{2}-3x-4}{2x^{3}+4x^{2}-6x+4}$ $\approx 6\cdot 10^{27}$ T 55 ms $38$ ms <1<1 ms $30$ ms
$\cfrac{2x+1}{x^{2}+x+1}$ $1926284$ A 88min 99s 1919 ms <1<1 ms $24$ ms
@@
Таблица 2: Output and computation time of our implementation of Algorithm 3 and timings of competing algorithms on a few examples. Here the output A stands for ‘‘Algebraic’’, and T stands for ‘‘Transcendental.’’
@@ 6.3.2 Large Random Inputs
Certifying algebraicity using our approach is difficult. Both the theoretical complexity and the timings of our algorithm suggest so, already for small degree and height of the input. Proving transcendence, however, is in general much easier as we expect that in this case – except for very specific polynomials – the $p$-curvatures for some small primes $p$ will not vanish. This might seem counterintuitive when compared to numbers – for which proving transcendence is considered hard – but this behavior was already observed in the functional context in [BSS25]. By ‘‘small’’, we mean that we expect that it suffices to check a number of $p$-curvatures significantly smaller than the bound $\sigma$ of Theorem 1.3, but the latter is, to our knowledge, the best proven bound. Thus, in practice, our algorithm should be viewed rather only as a semi-algorithm in the sense that if the output is algebraic the algorithm will not terminate in reasonable time, except in very specific examples. The observed behavior worsens as the input degree and height increase, hence let us focus on the case where the output is transcendental.
@@
Таблица 3: Average computation time of various algorithms on random rational function inputs of prescribed degree and height.
The timings are given for random inputs $a(x),b(x)\in\mathbb{Z}[x]$ of multiplicative height at most $H$ and degree n1n-1, respectively, $n$, where Algorithm 3 returns ‘‘Transcendental.’’ The tests for degrees up to $40$ were performed on samples of hundreds of random rational functions, the size of the sample for the tests in degree $80$ and $160$ were respectively $100$ and $10$. In all our experiments, we observed that most examples return transcendental after computing a non-vanishing $p$-curvature with $p$ no greater than $17$. In such cases, we avoid the computation of the Rothstein-Trager polynomial and only compute its leading coefficients, the discriminant of $b(x)$, whereas the factorization approach cannot take this shortcut.
@@
One can hand-craft examples on which we need to test relatively large primes $p$ – at least compared to the generic behavior. In the case of a quadratic polynomial, this smallest prime for which it does not factor into a product of two linear factors is the smallest quadratic non-residue in its splitting field. Examples of quadratic fields with prescribed least quadratic non-residue can be found in [MT21]. From this one can construct coefficients $u(x)$ of differential equations, for which the first prime for which the $p$-curvature does not vanish is comparatively large. For instance, if $a(x)=1$, $b(x)=x^{2}-3818929$, the prime $p=2$ divides $\Delta=\operatorname{res}(b(x),b^{\prime}(x))$, and for all other primes up to $43$ the $p$-curvatures vanish, but not the $47$-curvature. Our algorithm runs in $80$ milliseconds on this example.
@@
Список литературы
  • [AK19] Jeoung-Hwan Ahn and Soun-Hi Kwon ‘‘An Explicit Upper Bound for the Least Prime Ideal in the Chebotarev Density Theorem’’ In Annales de l’Institut Fourier 69.3, 2019, pp. 1411–1458 DOI: 10.5802/aif.3274
  • [Ax71] James Ax ‘‘On Schanuel’s conjectures’’ In Ann. of Math. (2) 93, 1971, pp. 252–268 DOI: 10.2307/1970774
  • [BCR24] Alin Bostan, Xavier Caruso and Julien Roques ‘‘Algebraic Solutions of Linear Differential Equations: An Arithmetic Approach’’ In Bulletin of the American Mathematical Society 61.4, 2024, pp. 609–658 DOI: 10.1090/bull/1835
  • [BCS15] Alin Bostan, Xavier Caruso and Йric Schost ‘‘A fast algorithm for computing the $p$-curvature’’ In ISSAC’15—Proceedings of the 2015 ACM International Symposium on Symbolic and Algebraic Computation ACM, New York, 2015, pp. 69–76
  • [BCS16] Alin Bostan, Xavier Caruso and Йric Schost ‘‘Computation of the similarity class of the $p$-curvature’’ In Proceedings of the 2016 ACM International Symposium on Symbolic and Algebraic Computation ACM, New York, 2016, pp. 111–118 DOI: 10.1145/2930889.2930897
  • [BH89] F. Beukers and G. Heckman ‘‘Monodromy for the Hypergeometric Function $_{n}F_{n-1}$’’ In Inventiones Mathematicae 95.2, 1989, pp. 325–354 DOI: 10.1007/BF01393900
  • [Bos+17] Alin Bostan et al. ‘‘Algorithmes Efficaces en Calcul Formel’’ 686 pages. Imprimй par CreateSpace. Aussi disponible en version йlectronique Palaiseau: Frйdйric Chyzak (auto-йdit.), 2017 URL: https://hal.archives-ouvertes.fr/AECF/
  • [Bou97] Auguste Boulanger ‘‘Contribution а l’йtude des йquations diffйrentielles linйaires et homogиnes intйgrables algйbriquement’’, 1897
  • [BS09] Alin Bostan and Йric Schost ‘‘Fast algorithms for differential equations in positive characteristic’’ In Proceedings of the 2009 international symposium on symbolic and algebraic computation, ISSAC 2009, Seoul, July 28–31, 2009. New York, NY: Association for Computing Machinery (ACM), 2009, pp. 47–54 DOI: 10.1145/1576702.1576712
  • [BSS25] Alin Bostan, Bruno Salvy and Michael F. Singer ‘‘On Deciding Transcendence of Power Series’’, 2025 DOI: 10.48550/arXiv.2504.16697
  • [CC85] D.. Chudnovsky and G.. Chudnovsky ‘‘Applications of Padй approximations to the Grothendieck conjecture on linear differential equations’’ In Number theory (New York, 1983–84) 1135, Lecture Notes in Math. Springer, Berlin, 1985, pp. 52–100 DOI: 10.1007/BFb0074601
  • [Chu83] G.. Chudnovsky ‘‘On the method of Thue-Siegel’’ In Ann. Math. (2) 117, 1983, pp. 325–382 DOI: 10.2307/2007080
  • [Fid85] Charles M. Fiduccia ‘‘An efficient formula for linear recurrences’’ In SIAM J. Comput. 14.1, 1985, pp. 106–112 DOI: 10.1137/0214007
  • [FY24] Florian Fьrnsinn and Sergey Yurkevich ‘‘Algebraicity of Hypergeometric Functions with Arbitrary Parameters’’ In Bulletin of the London Mathematical Society, 2024, pp. 23 pp. DOI: 10.1112/blms.13103
  • [Her73] Ch. Hermite ‘‘Sur la fonction exponentielle’’ In C. R. Acad. Sci., Paris 77, 1873, pp. 18–24
  • [Her93] Ch. Hermite ‘‘Sur la gйnйralisation des fractions continues algйbriques, Extrait d’une lettre а M. Pincherlй’’ In Annali di Mat. (2) 21, 1893, pp. 289–308 DOI: 10.1007/BF02420446
  • [Hon81] Taira Honda ‘‘Algebraic Differential Equations’’ In Symposia Mathematica XXIV (Sympos., INDAM, Rome, 1979) London-New York: Academic Press, 1981, pp. 169–204
  • [Jac37] Nathan Jacobson ‘‘Abstract Derivation and Lie Algebras’’ In Transactions of the American Mathematical Society 42.2, 1937, pp. 206 DOI: 10.2307/1989656
[…]
  • [Jag64a] H. Jager ‘‘A multidimensional generalization of the Padй table. IV, V, VI’’ In Nederl. Akad. Wet., Proc., Ser. A 67, 1964, pp. 227–249
  • [Kat72] Nicholas M. Katz ‘‘Algebraic Solutions of Differential Equations ($p$-Curvature and the Hodge Filtration)’’ In Inventiones Mathematicae 18.1-2, 1972, pp. 1–118 DOI: 10.1007/BF01389714
  • [Kat82] Nicholas M. Katz ‘‘A Conjecture in the Arithmetic Theory of Differential Equations’’ In Bulletin de la Sociйtй mathйmatique de France 79, 1982, pp. 203–239 DOI: 10.24033/bsmf.1960
  • [Kro80] Leopold Kronecker ‘‘Ьber die Irreducibilitдt von Gleichungen’’ In Monatsberichte der Kцniglichen Preussische Akademie des Wissenschaften zu Berlin, 1880, pp. 155–163
  • [KW22] Habiba Kadiri and Peng-Jie Wong ‘‘Primes in the Chebotarev Density Theorem for All Number Fields (with an Appendix by Andrew Fiori)’’ In Journal of Number Theory 241, 2022, pp. 700–737 DOI: 10.1016/j.jnt.2022.03.012
  • [Li12] Xiannan Li ‘‘The Smallest Prime That Does Not Split Completely in a Number Field’’ In Algebra & Number Theory 6.6, 2012, pp. 1061–1096 DOI: 10.2140/ant.2012.6.1061
  • [LMO79] J.. Lagarias, H.. Montgomery and A.. Odlyzko ‘‘A Bound for the Least Prime Ideal in the Chebotarev Density Theorem’’ In Inventiones Mathematicae 54.3, 1979, pp. 271–296 DOI: 10.1007/BF01390234
  • [Loo83] Rьdiger Loos ‘‘Computing rational zeros of integral polynomials by $p$-adic expansion’’ In SIAM J. Comput. 12.2, 1983, pp. 286–293 DOI: 10.1137/0212017
  • [Mah68] K. Mahler ‘‘Perfect systems’’ In Compos. Math. 19, 1968, pp. 95–166 URL: https://eudml.org/doc/88959
  • [MT21] Kevin McGown and Enrique Treviсo ‘‘The least quadratic non-residue’’ In Mexican mathematicians in the world—trends and recent contributions 775, Contemp. Math. Amer. Math. Soc., 2021, pp. 205–231 DOI: 10.1090/conm/775/15593
  • [Mur94] V. Murty ‘‘The Least Prime Which Does Not Split Completely’’ In Forum Mathematicum 6.6, 1994 DOI: 10.1515/form.1994.6.555
  • [Pad01] H. Padй ‘‘Sur l’expression gйnйrale de la fraction rationnelle approchйe de $(1+x)^{m}$.’’ In C. R. Acad. Sci., Paris 132, 1901, pp. 754–756
  • [Pad92] H. Padй ‘‘Sur la reprйsentation approchйe d’une fonction par des fractions rationnelles’’ In Ann. Sci. Йc. Norm. Supйr. (3) 9, 1892, pp. 3–93 DOI: 10.24033/asens.378
  • [Pad94] H. Padй ‘‘Sur la gйnйralisation des fractions continues algйbriques’’ In C. R. Acad. Sci., Paris 118, 1894, pp. 848–850
  • [Pag21] Raphaлl Pagиs ‘‘Computing Characteristic Polynomials of $p$-Curvatures in Average Polynomial Time’’ In Proceedings of the 2021 International Symposium on Symbolic and Algebraic Computation Virtual Event Russian Federation: ACM, 2021, pp. 329–336 DOI: 10.1145/3452143.3465524
  • [Pai87] Paul Painlevй ‘‘Sur es йquations diffйrentielles linйaires’’ In Comptes Rendus de l’Acadйmie des Sciences, Paris 105, 1887, pp. 1829–1832
  • [Put96] Marius Put ‘‘Reduction modulo $p$ of differential equations’’ In Indagationes Mathematicae 7.3, 1996, pp. 367–387 DOI: https://doi.org/10.1016/0019-3577(96)83726-8
  • [Ris70] Robert H. Risch ‘‘The Solution of the Problem of Integration in Finite Terms’’ In Bulletin of the American Mathematical Society 76.3, 1970, pp. 605–608 DOI: 10.1090/S0002-9904-1970-12454-5
  • [Rob55] Herbert Robbins ‘‘A Remark on Stirling’s Formula’’ In The American Mathematical Monthly 62.1 [Taylor Francis, Ltd., Mathematical Association of America], 1955, pp. 26–29 URL: http://www.jstor.org/stable/2308012
  • [Rot76] Michael Rothstein ‘‘Aspects of Symbolic Integration and Simplification of Exponential and Primitive Functions’’, 1976
  • [Rot77] Michael Rothstein ‘‘A New Algorithm for the Integration of Exponential and Logarithmic Functions’’ In Proceedings of the 1977 MACSYMA Users’ Conference, 1977
  • [RS62] J. Rosser and Lowell Schoenfeld ‘‘Approximate formulas for some functions of prime numbers’’ In Illinois Journal of Mathematics 6.1 Duke University Press, 1962, pp. 64–94 DOI: 10.1215/ijm/1255631807
  • [Sch82] Arnold Schцnhage ‘‘The fundamental theorem of algebra in terms of computational complexity’’ In Manuscript. Univ. of Tьbingen, Germany, 1982
  • [Sin79] Michael F. Singer ‘‘Algebraic Solutions of $n$th Order Linear Differential Equations’’ In Proceedings of the Queen’s University 1979 Conference on Number Theory 54, Queens Papers in Pure and Applied Mathematics, 1979, pp. 379–420
  • [Sta80] R.P. Stanley ‘‘Differentiably Finite Power Series’’ In European Journal of Combinatorics 1.2, 1980, pp. 175–188 DOI: 10.1016/S0195-6698(80)80051-5
  • [SZ94] Bruno Salvy and Paul Zimmermann ‘‘GFUN: A Maple Package for the Manipulation of Generating and Holonomic Functions in One Variable’’ In ACM Transactions on Mathematical Software 20.2, 1994, pp. 163–177 DOI: 10.1145/178365.178368
  • [Tra76] Barry M. Trager ‘‘Algebraic Factoring and Rational Function Integration’’ In Proceedings of the Third ACM Symposium on Symbolic and Algebraic Computation - SYMSAC ’76 Yorktown Heights, New York, United States: ACM Press, 1976, pp. 219–226 DOI: 10.1145/800205.806338
  • [Tsc26] N. Tschebotareff ‘‘Die Bestimmung der Dichtigkeit einer Menge von Primzahlen, welche zu einer gegebenen Substitutionsklasse gehцren’’ In Mathematische Annalen 95.1 Springer Science and Business Media LLC, 1926, pp. 191–228 DOI: 10.1007/bf01206606
  • [VV00] Jeffrey D Vaaler and Josй Felipe Voloch ‘‘The Least Nonsplit Prime in Galois Extensions of $\mathbb{Q}$’’ In Journal of Number Theory 85.2, 2000, pp. 320–335 DOI: 10.1006/jnth.2000.2551
  • [vzGG13] Joachim Gathen and Jьrgen Gerhard ‘‘Modern Computer Algebra’’ Cambridge: Cambridge university press, 2013
  • [Чеб23] Н Чеботарев ‘‘Определение плотности совокупности простых чисел, принадлежащих к заданному классу подстановок. I’’ In Известия Российской Академии Наук 17.1-18, VI, 1923, pp. 205–230
  • [Чеб23a] Н Чеботарев ‘‘Определение плотности совокупности простых чисел, принадлежащих к заданному классу подстановок. II’’ In Известия Российской Академии Наук 17.1-18, VI, 1923, pp. 231–250
Faculty of Mathematics, University of Vienna, Oskar-Morgenstern-Platz 1, 1090, Vienna, Austria
CNRS; Laboratoire de Mathйmatiques de Versailles, UMR-8100, UVSQ, 45 avenue des Йtats-Unis 78000, Versailles CEDEX, France

### v2

@@ 5.5 Complexity Estimate
Proposition 5.15.
We have
$\delta(\Delta_{R})<\frac{4}{(\log 2)^{2}}\left((2n-1)^{2}\log(H)+4n^{2}\log(n+1)+\log\left(\frac{2^{4n^{2}-1}}{3^{n/2}(\pi n)^{2n-1}}\right)\right)^{2}.$ (17)
In particular, with the notations as above, we have $\delta(\Delta_{R})=\tilde{O}(n^{4}(\log H+\log n)^{2})$.
Proof.
The statement can be deduced by applying a bound on the size of coefficients of the factors of the Rothstein-Trager resultant, Proposition 5.7, and then Proposition 5.3 to compute a bound on $\Delta_{R}$: first we compute a bound on the height of $R(w)$, the square-free part of the Rothstein-Trager resultant of $a(x)$ and $b(x)$, and use a bound on the height of its Rothstein-Trager to obtain a bound for its leading coefficient, which is precisely $\Delta_{R}$. We conclude using Lemma 5.16, below. ∎
@@
Let $\Delta\in\mathbb{N}$ and $\delta=\prod_{p|\Delta}p^{1/(p-1)}$, then $\delta(\Delta)\leq\frac{4}{\log(2)^{2}}\log(\Delta)^{2}$.
Proof.
First we remark that if $\ell\in\mathbb{N}$ is such that $\ell\#\leq\Delta<(\ell+1)\#$, then $\delta(\Delta)\leq\delta(\ell\#)$ where $\ell\#\coloneqq\prod_{p\leq\ell}p$ is the primorial of $\ell$. This follows from the fact that $x^{1/(x-1)}$ is decreasing for $x\geq 2$. We will now estimate $\delta(\ell\#)$. In [41, Theorem 10], we find that for $\ell\geq 557$, we have $0.92\cdot\ell\leq\log(\ell\#)$. Checking numerically that for $1\leq\ell\leq 557$ we have $2^{\ell/2}\leq\ell\#$ we conclude that this inequality holds for all $\ell\geq 1$. Moreover we have the estimate $\sum_{p\leq\ell}\frac{\log(p)}{p-1}\leq 2\sum_{p\leq\ell}\frac{\log(p)}{p}\leq 2\log(\ell)$ following [41, Eq. (2.5), (2.11)]. Taking now $\ell$ such that $\ell\#\leq\Delta\leq(\ell+1)\#$ we obtain that $\delta(\Delta)\leq\frac{4}{\log(2)^{2}}\log(\Delta)^{2}$. ∎
@@
Given $u(x)\coloneqq\frac{a(x)}{b(x)}\in\mathbb{Q}(x)$ of degree $n$ and multiplicative height $H$, deciding if all solutions of $y^{\prime}(x)=u(x)y(x)$ are algebraic can be done using Algorithm 3 by performing $\tilde{O}(\Delta_{b}^{6}B)=\tilde{O}(H^{12n-6}n^{12n}3^{-3n})$ bit operations, where $\tilde{O}$ hides factors polynomial in $n$ and logarithmic in $H$.
Proof.
We compute $R(w),\Delta_{b},$ and a bound on $\delta$ as described in the algorithm, and the maximal modulus of a root of $R(w)$, up to a constant factor by for instance setting $\tau=\frac{1}{2}$ in Proposition 5.12. This we will use as an upper bound on $B$. With this we obtain an upper bound on $\sigma$. Clearly the complexities of these computations are negligible compared to the computation of sufficiently many $p$-curvatures as described below.
By Theorem 1.3, it suffices to compute the $p$-curvatures up to $\sigma$, defined in its statement. Thus, we want to set $S=\sigma$ in Proposition 5.13. In accordance with Proposition 5.15 the contribution of $\delta$ in the size of $\sigma$ is negligible and we have $\sigma=\tilde{O}(B\Delta_{b}^{6})$, where we omit logarithmic factors in $\Delta_{b}$. Cauchy’s bound on the roots of $R(w)=r_{n}w^{n}+\dots+r_{0}$ states that we can take $B$ at most $1+\max_{i}\left|\frac{r_{i}}{r_{n}}\right|$, hence $\sigma=\tilde{O}(\Delta_{b}^{5}H_{R})$, where $H_{R}$ denotes the multiplicative height of $R(w)$. Using estimates from Proposition 5.3 to estimate this height and $\Delta_{b}$ we find both $\Delta_{b}$ and $H_{R}$ to be in $\tilde{O}(H^{2n-1}n^{2n-1/2}3^{-n/2})$. Taking the $6$-th power of this quantity and combining with Proposition 5.13 concludes. ∎
@@ 6.1 Algorithmic Improvements
The crucial algorithmic idea of the design of Algorithm 3 is to delay the moment when costly computations are done, that is the computation of the Rothstein-Trager resultant $R(w)$, when the degree of the input increases. By checking the ‘‘first’’ $p$-curvatures, in our implementation for primes $p$ up to $\Delta_{b}$, we try to return ‘‘Transcendental’’ before having to compute $R(w)$. In case no $p$-curvature vanishes until this point, the computation of further $p$-curvatures, potentially up to $\sigma$, is inevitable and this has much higher complexity than the computation of $R(w)$. When computing the bound $B$ on the modulus of the roots of $R(w)$, the precision $\tau$ of the computation influences the complexity, see Proposition 5.12. In our case, computing all $p$-curvatures up to $\sigma$ already exceeds the capabilities of modern computers in most cases. For theoretical estimates, having $B$ multiplied by a constant factor does not change the complexity of the computations. Our implementation relies on native SageMath functions numerically computing zeroes of a polynomial, up to a fixed precision of $2^{-53}$.
The naïve computation of $\delta(\Delta_{R})$ can be costly, as it relies on the factoring of large integers. Using the bound obtained in (17) we avoid the computation of $\Delta_{R}$, an iterated resultant, and its prime factors. However, this bound appears to be highly non-optimal on examples. In our implementation we thus decide to compute $\Delta_{R}$ and try for a small number of primes (up to $T=1009$) to find its contributions to $\delta(\Delta_{R})$. There remains $\tilde{\Delta}_{R}$, the product of its large factors, whose contribution to $\delta(\Delta_{R})$ is bounded from above by $T^{v/(T-1)}$ for $v=\log_{T}(\tilde{\Delta}_{R})$.
@@ 6.2.1 Computation of $p$-curvatures
First of all we would like to ensure and to convince the reader that our implementation of Algorithm 4 of [9] to compute $p$-curvatures of first-order differential equations is indeed faster than Pagиs’ implementation of his algorithm [37] applied to order one equations. The following timings in average were noted for both algorithms when asked to compute $p$-curvatures for primes up to $S$ on random polynomial inputs $a(x),b(x)\in\mathbb{Z}[x]$ of given multiplicative height $H$ and degrees n1n-1, respectively $n$.
@@ 33 $2^{80}$ $100$ $25$ ms $140$ ms
Table 1: Computation time of two algorithms computing $p$-curvatures.
@@ 6.2.2 Maple’s istranscendental
The maple package gfun [44] has a command called istranscendental based on the algorithms described in [8], aiming to prove transcendence of a given solution of a given differential equation. Among other things, it checks whether the singularities of the minimal differential operator annihilating a given function are all regular, and if its local exponents are rational. This is a necessary criterion for the algebraicity of the solution, so in case one of these conditions is violated, the command outputs true. For order one differential equations however, Proposition 3.3 ensures that the conditions are also sufficient for algebraicity. While istranscendental outputs FAIL in this case, we can actually conclude algebraicity and thus we are (ab)using the command as an implementation of Algorithm 2.
@@ 6.3.1 Small Inputs
$\cfrac{3x-4}{2x^{2}-6x+4}$ $265$ A $156$ ms $45$ ms <1<1 ms $25$ ms
$\cfrac{7x^{2}-3x-4}{2x^{3}+4x^{2}-6x+4}$ $\approx 3\cdot 10^{29}$ T 55 ms $38$ ms <1<1 ms $30$ ms
$\cfrac{2x+1}{x^{2}+x+1}$ $3851774$ A 1414min $44$s 1919 ms <1<1 ms $24$ ms
@@
Table 2: Output and computation time of our implementation of Algorithm 3 and timings of competing algorithms on a few examples. Here the output A stands for ‘‘Algebraic’’, and T stands for ‘‘Transcendental.’’
@@ 6.3.2 Large Random Inputs
Certifying algebraicity using our approach is difficult. Both the theoretical complexity and the timings of our algorithm suggest so, already for small degree and height of the input. Proving transcendence, however, is in general much easier as we expect that in this case – except for very specific polynomials – the $p$-curvatures for some small primes $p$ will not vanish. This might seem counterintuitive when compared to numbers – for which proving transcendence is considered hard – but this behavior was already observed in the functional context in [8]. By ‘‘small’’, we mean that we expect that it suffices to check a number of $p$-curvatures significantly smaller than the bound $\sigma$ of Theorem 1.3, but the latter is, to our knowledge, the best proven bound. Thus, in practice, our algorithm should be viewed rather only as a semi-algorithm in the sense that if the output is algebraic the algorithm will not terminate in reasonable time, except in very specific examples. The observed behavior worsens as the input degree and height increase, hence let us focus on the case where the output is transcendental.
@@
Table 3: Average computation time of various algorithms on random rational function inputs of prescribed degree and height.
The timings are given for random inputs $a(x),b(x)\in\mathbb{Z}[x]$ of multiplicative height at most $H$ and degree n1n-1, respectively, $n$, where Algorithm 3 returns ‘‘Transcendental.’’ The tests for degrees up to $40$ were performed on samples of hundreds of random rational functions, the size of the sample for the tests in degree $80$ and $160$ were respectively $100$ and $10$. In all our experiments, we observed that most examples return transcendental after computing a non-vanishing $p$-curvature with $p$ no greater than $17$. In such cases, we avoid the computation of the Rothstein-Trager polynomial and only compute its leading coefficients, $\operatorname{res}_{x}(b(x),b^{\prime}(x))$, whereas the factorization approach cannot take this shortcut.
@@
One can hand-craft examples on which we need to test relatively large primes $p$ – at least compared to the generic behavior. In the case of a quadratic polynomial, this smallest prime for which it does not factor into a product of two linear factors is the smallest quadratic non-residue in its splitting field. Examples of quadratic fields with prescribed least quadratic non-residue can be found in [30]. From this one can construct coefficients $u(x)$ of differential equations, for which the first prime for which the $p$-curvature does not vanish is comparatively large. For instance, if $a(x)=1$, $b(x)=x^{2}-3818929$, the prime $p=2$ divides $\Delta_{b}=\operatorname{res}(b(x),b^{\prime}(x))$, and for all other primes up to $43$ the $p$-curvatures vanish, but not the $47$-curvature. Our algorithm runs in $80$ milliseconds on this example.